OpenStudy (anonymous):

i need help solving the problem (1/A)-(1/B)=(1/C) and i need to solve for A

OpenStudy (anonymous):

bc/(b+c)

OpenStudy (anonymous):

exactly.. then you just solve for A

OpenStudy (anonymous):

so do you multipb the b and the c together so the denomonators are the same that way you can divide like that?

OpenStudy (anonymous):

yes, you want to solve for A you can find that [A,B]=AB and transform the denominators into AB by multiplying by A and B..

OpenStudy (anonymous):

so do i subtract (1/A) the retricethe (1/C) so that my formula is (1/B)-(1/C)=(1/A) and then make is so that (C/CB)-(B/BC)=(1/A)

OpenStudy (anonymous):

that last part is where i am getting confused

OpenStudy (anonymous):

how are you getting BC into the numerator?

OpenStudy (anonymous):

do you know what less common multiple of two numbers is?

OpenStudy (anonymous):

*least

OpenStudy (anonymous):

yeah i understand that part but the letters are what is confusing me

OpenStudy (anonymous):

we have that LCM{X,Y}=X*Y/GCD{X,Y} for the needs of the exercise we assume thatGCD{X,Y}=1 and we have LCM{X,Y}=X*Y... that's all

OpenStudy (anonymous):

i fluttering give up. i've wasted to much of my time and my parents money to care anymore...

OpenStudy (anonymous):

as you wish..

OpenStudy (anonymous):

sorry i just dont get it, thats okay thanks for your help anyways! i really do appreiate it!

OpenStudy (anonymous):

ok..